# From CVE Entries to Verifiable Exploits: An Automated Multi-Agent Framework for Reproducing CVEs

### v1

@@ 1 Introduction
Tens of thousands of software vulnerabilities are discovered every year, with more than 49,000 vulnerabilities tracked by the National Vulnerability Database (NVD) [42] in 2024 alone. Many of these vulnerabilities are cataloged using the CVE format, which provides a standardized identification system to help organizations address security flaws in both software and hardware. Still, existing vulnerability detection tools, such as those listed by OWASP [44], struggle to keep up with the increasing volume of code released each year. As a result, many vulnerabilities remain undiscovered in codebases for years (e.g., up to 2 years in Chromium and 7 years in OpenSSL [3]) posing significant security risks.
A key requirement for developing effective automated vulnerability detection techniques is the availability of high-quality datasets with detailed examples and ground truth. Studies have shown that low-quality data can cause machine learning (ML) based vulnerability detection techniques to learn spurious correlations [4, 49]. Although public repositories, such as the NVD, include hundreds of thousands of CVEs, they often lack critical details, such as the vulnerable code, setup instructions for the vulnerable environment, working exploits, and validation steps, which are all required for the effective replication of a vulnerability and to build suitable ground truth datasets for research [38]. This is because the main purpose of vulnerability advisories is to alert system administrators about software that needs to be updated, and concrete exploit code is often kept confidential for ethical reasons. Unfortunately, generating this missing information from CVEs alone requires both substantial manual effort and extensive security expertise. For instance, Mu et al. [38] spent over 3,600 hours of manual effort with 43 security experts to reproduce publicly reported CVEs but managed to reproduce only 368 memory-related vulnerabilities in Linux.
@@
To address these challenges, the security community has adopted several approaches: manual annotation [21, 65], inserting bugs into real-world code [41, 36], and heuristic-based mining [13, 10, 40, 15, 7]. While these methods have advanced the field, they come with limitations, such as limited support for programming languages and Common Weakness Enumerations (CWE) categories, mislabeled vulnerabilities [50], and data becoming stale. For instance, the DARPA Cyber Grand Challenge dataset from 2016 [11] quickly became outdated due to its small size and low-complexity bugs [20]. On the other hand, automated data collection helps address some of these limitations, but often tend to be limited in scope. For instance, ARVO [34] automatically curates reproducible builds with triggering inputs from OSS-Fuzz reports, but its scope is limited to memory corruption bugs in C/C++. This highlights a clear need for a dataset curation method that enables fully reproducible vulnerability instances across a wide range of programming languages and vulnerability types.
Recently, LLMs have demonstrated remarkable capabilities in solving complex software engineering (SWE) and security tasks, ranging from automated bug detection [39, 16] and repair [46, 57, 24], to writing fuzzer test harnesses [12]. Moreover, the shift to agentic workflows [63, 8, 62, 32], in which LLM-powered agents are provided access to real-world tools [1], has further pushed the capabilities of general purpose LLMs and lead to significant performance gains on extremely difficult real-world SWE related benchmarks, such as SWE-bench [25]. This raises the question whether LLM agents could also be useful in automatically setting up and reproducing existing CVEs, allowing us to build reliable, large, and realistic benchmarks for the security community.
In this paper, we present CVE-Genie, a fully-automated, LLM-driven multi-agent framework for end-to-end CVE reproduction, which embodies what we define as key properties of an ideal CVE reproduction system, captured by the acronym EAGER. Specifically, an ideal system: Generates working Exploit/proof-of-concept (PoC); Builds Assessors/verifiers for the exploit; Generalizes across CWEs, languages, and projects; Is End-to-end automated; and Rebuilds vulnerable environments. This paper makes the following contributions:
@@
    We develop CVE-Genie, a four-stage, end-to-end automated pipeline; Processor, Builder, Exploiter, and CTF Verifier, that extracts CVE data (i.e., source code, security advisories, patches, etc.), rebuilds the environment, generates exploits and verifiers to assess the exploits, and validates the final results.
@@
    We systematically select the optimal configuration of LLMs for CVE-Genie for optimal performance over diverse CVEs, encompassing diverse range of vulnerabilities, programming languages, and projects, including scenarios with incomplete CVE information.
@@
    We ran CVE-Genie on 841 CVEs (published between June 2024–May 2025), successfully reproducing 428 CVEs across 267 projects, 141 CWEs, and 22 programming languages, highlighting our generalizability.
@@
    We systematically evaluate CVE-Genie’s design and demonstrate that each component is essential for optimal performance: even advanced LLMs like o3 fail to reproduce even one CVE without our agentic framework.
@@
    We open-source our framework, source code, and datasets of reproduced CVEs as well as logs for all agents conversations, providing an ongoing valuable contribution to the community for downstream research tasks (e.g., tools for automated vulnerability detection).
@@ 2 Background and Related Work
Existing Benchmarks and Limitations. To evaluate various vulnerability detection techniques (rule-based methods [44, 58], ML-approaches [36, 30, 31, 29], and LLM-based systems [17, 19, 47, 18]), the community needs standardized benchmarks. Prior work has explored four main approaches for addressing these challenges. (1) Manual efforts involve analyzing [21] or reproducing [66] vulnerabilities in real-world code, or crafting synthetic examples [54], but datasets are small, lack ecological validity, and often suffer labeling errors due to human mistakes (e.g., 50% data in Devign [65] have wrong labels  [50]). (2) Automated labeling using static analysis (SA) tools, such as bandit [6] or infer [23], reduces manual effort but introduces many false positives [48, 64]. (3) Mining real-world data, such as GitHub issues or CVE patch commits, assumes that all removed or modified code is vulnerable; However, this assumption often leads to mislabeled samples, as Risse et al. [50] highlighted for DiverseVul [10] and BigVul [15]. (4) Relying on proven sources, such as OSS-Fuzz, offers verified exploits and sanitizer feedback; However, these are restricted to certain bug types, primarily memory corruption in C/C++, and do not generalize across languages or vulnerability classes [34].
@@
Call for EAGER Style for CVE Reproduction. A CVE ID is a unique identifier for publicly disclosed security vulnerabilities in real-world software. Each CVE ID corresponds to a CVE entry that includes key details for that CVE, such as vulnerability description, source code, affected versions, security advisories, CWE classifications, and patches, enabling security professionals to assess and mitigate risks. Creating a dataset from the standard CVE database is a promising way towards addressing the aforementioned challenges (see Appendix B). Therefore, we introduce EAGER, a set of crucial criteria for ideal CVE reproduction:
@@
However, none of existing benchmarks satisfy all the criteria (Table 2). For example, manual effort for producing EAGER datasets does not scale and is not generalizable [38, 66]. ARVO [34] focuses solely on memory-related bugs from OSS-Fuzz C/C++ projects, limiting its scope. In contrast, CVE-Genie is the only framework satisfying all criteria.
@@
Large Language Models and Agents. LLMs are transformer-based neural network models with billions of parameters. They have shown remarkable performances in various application domains, such as question answering [43], mathematical reasoning [2], and code generation [9, 5]. Recent work also shows their utility in security tasks, e.g., such as bug detection and repair [46, 57, 24] and writing fuzzer test harnesses [12].
LLM agents are systems that combine LLMs with tools. They can finish complex tasks via a sequence of actions, including planning, generation, and tool executions [53, 22, 45, 27, 26]. An agentic system can have multiple agents, each with a sub-task, such as reasoning, critiquing, generating, perceiving, remembering, or teaching, as well as its own workflow and tool sets. In security, researchers have started to construct LLM agents to find and repair vulnerabilities. There are multi-agent patching systems, where different sub-agents are responsible for fault localization, patch generation and validation (e.g., PatchPilot [28], and PatchAgent [61]). Through the recent DAPRA AIxCC competition, researchers constructed agentic-based systems with end-to-end capabilities for vulnerability detection and patching [18]. These recent successes provides the argument for using LLM agents in vulnerability analysis, software development, and external tool interaction, key capabilities for building an EAGER-style automated end-to-end CVE reproduction framework.
@@ 3 CVE-Genie
Like human analysts [38], CVE-Genie utilizes four core modules to process and reproduce CVEs: (1) the Processor (Section 3.1), which retrieves source code and constructs a structured knowledge base; (2) the Builder (Section 3.2), which reconstructs the vulnerable environment; (3) the Exploiter (Section 3.3), which reproduces the exploit; and (4) the CTF Verifier (Section 3.4), which generates a verifier for the exploit. We also design and integrate a set of tools that enable LLM agents in these modules to perform their respective tasks (see Appendix A). To illustrate the framework in practice, we also present a full end-to-end reproduction of CVE-2024-4340 by CVE-Genie (Figure 2).
To bridge the gap identified in Section 2, namely the lack of any system capable of true EAGER-style CVE reproduction, CVE-Genie’s design is guided by the following principles.
@@
    Modular Task Decomposition: Since LLMs struggle with long, complex contexts [54] and tasks [25], CVE-Genie decomposes reproduction into focused modules handled by specialized agents with systematically engineered prompts [^1] All prompts will be provided in the open-source artifact..
@@
    Robustness to Incomplete Data: Given that missing CVE fields can reduce reproduction success by up to 44% [38], agents in CVE-Genie can fall back on patch analysis and direct source-code inspection when advisories or PoCs are unavailable.
@@
    Reliability through Self-Critique: To counter LLMs’ reasoning limits [54], each module uses paired developer and critic agents in a ReAct-style loop [60], enabling iterative refinement and internal verification.
Together, these principles ensure CVE-Genie can deliver the first practical, automated, and generalizable framework for EAGER-style CVE reproduction. The following subsections detail the workings of each CVE-Genie component, using CVE-2024-4340 as a running example to illustrate the end-to-end reproduction shown in Figure 2.
Figure 2: CVE-Genie architecture and an end-to-end example of workflow of reproduction for CVE-2024-4340, i.e., Denial of Service due to RecursionError in sqlparse < v0.5.0.
@@ 3.1 Processor
This module comprises of two sub-modules: the Data Processor, which locates the vulnerable project’s source code and gathers relevant raw resources, and the LLM-based Knowledge Builder, which transforms these resources into a structured knowledge base. Below we provide detailed functionalities of these sub-modules.
@@ 3.1.1 Data Processor 1
Goal/Rationale. For the given CVE ID, this sub-module identifies and extracts the vulnerable version of the software along with all relevant CVE details, as follows:
Source Code Extraction. The Data Processor begins by locating the source code for the vulnerable project linked to the CVE, an essential requirement for CVE-Genie to reproduce the vulnerability. It searches the GitHub URLs referenced in the “cvelist”222cvelist” is an automated pilot program for CVE submission through GitHub (https://github.com/CVEProject/cvelist) to identify the project repository. For closed-source software, Data Processor provides the option to the user to supply the repository URL. Once the source code is located, it identifies the affected software configurations: specific versions, platforms, or settings vulnerable to the CVE. For instance, CVE-2024-4340 affects all versions of “sqlparse” prior to v0.5.0. Based on this, the latest affected (vulnerable) version (v0.4.4) is retrieved using the GitHub API, and its source code is downloaded (as shown in Figure 2).
Vulnerability Information Extraction. After the vulnerable version of the project source code is downloaded, Data Processor extracts the following four key pieces of information from “cvelist”, if available (see Figure 2):
  1. 1.
    CVE description: A high-level summary of the vulnerability in the target project.
  2. 2.
    CWE data: The CWE information, a categorization system for hardware and software weaknesses that associates the given CVE with a speficic type of vulnerability, e.g., CWE-674 (Uncontrolled Recursion) [^3] https://cwe.mitre.org/data/definitions/674.html.
  3. 3.
    Patch commits: The git-diff of code changes in the target source code that fix the vulnerability. These commits help LLMs localize the issue, understand its root cause [21, 13, 15], and generate effective exploits.
  4. 4.
    Security advisories and PoC: We filter URLs referenced in the CVE entry by keywords (e.g., “security”, “advisory”, “advisories”, “bounty”, “bounties”), and scrape their contents.
@@ 3.1.2 Knowledge Builder 2
Goal/Rationale. This agent is instructed to analyze, distill, and organize the raw information of, extracted by the Data Processor, into a structured knowledge base while retaining essential details required to reproduce the given CVE. This includes extracting PoC details or step-by-step instructions from security advisories, which are invaluable for accurately recreating the vulnerability. For instance, CVE-2024-4340 has a PoC provided in its security advisory and the patch commit helps understand the root cause by highlighting where the vulnerability was mitigated. As shown in Figure 2, the Knowledge Builder included both of these details in the knowledge base. In case an advisory does not include an exploit, the LLM agent is tasked with generating an overview of what an exploit might entail. Additionally, capturing the patch details helps to localize the root cause of the vulnerability [21, 13, 7] and aids in crafting an effective exploit. This knowledge base acts as the long-term memory for agents, helping them to effectively reproduce CVEs. By including only essential information, the knowledge base avoids overloading the LLM’s context window, ensuring efficient reproduction.
Output. A structured CVE knowledge base providing details of CVE including its associated CWEs, affected software configurations, root causes (if provided in the security advisory), and essential information from security advisories and patch commits necessary for exploitation.

### v2

@@ 1 Introduction
Tens of thousands of software vulnerabilities are discovered every year, with more than 40,000 vulnerabilities tracked by the National Vulnerability Database (NVD) [42] in 2025 alone. Many of these vulnerabilities are cataloged using the CVE format, which provides a standardized identification system to help organizations address security flaws in both software and hardware. Still, existing vulnerability detection tools, such as those listed by OWASP [44], struggle to keep up with the increasing volume of code released each year. As a result, many vulnerabilities remain undiscovered in codebases for years (e.g., up to 2 years in Chromium and 7 years in OpenSSL [3]) posing significant security risks.
Developing effective automated vulnerability detection techniques depends on high-quality datasets with reliable ground truth, as low-quality data can lead to unreliable evaluations [51] and to learn spurious correlations, especially in machine learning (ML) based approaches [4, 49]. Although public repositories, such as the NVD, include hundreds of thousands of CVEs, they often lack critical details, including vulnerable code, environment setup, working exploits, and validation steps, which are all required to effectively reproduce a vulnerability and to build ground truth datasets for research [38]. This is because the main purpose of vulnerability advisories is to alert system administrators about software that needs to be updated, and concrete exploit code is often kept confidential for ethical reasons. Unfortunately, generating this missing information from CVEs alone is labor-intensive and requires significant expertise, e.g., Mu et al. [38] spent over 3,600 hours with 43 security expertise to reproduce only 368 memory vulnerabilities in Linux.
@@
To address these challenges, the security community has adopted several approaches: manual annotation [21, 65], inserting bugs into real-world code [41, 36], and heuristic-based mining [13, 10, 40, 15, 7]. While these methods have advanced the field, they come with limitations, such as limited support for software and Common Weakness Enumerations (CWE) categories, mislabeled vulnerabilities [50], and data becoming stale. For instance, the DARPA Cyber Grand Challenge dataset from 2016 [11] quickly became outdated due to its small size and low-complexity bugs [20]. Meanwhile, automated data collection helps address some of these limitations, but often tends to be limited in scope, e.g., ARVO [34] automatically curates reproducible builds with triggering inputs from OSS-Fuzz reports, but its scope is limited to memory corruption bugs in C/C++. This highlights a clear need for a dataset curation method that enables fully reproducible vulnerability instances across a wide range of software and vulnerability types.
Recently, LLMs have demonstrated remarkable capabilities in solving complex software engineering (SWE) and security tasks, ranging from automated bug detection [39, 16] and repair [46, 57, 24], to writing fuzzer test harnesses [12]. Moreover, the shift to agentic workflows [63, 8, 62, 32], in which LLM-powered agents are provided access to real-world tools [1], has further pushed the capabilities of general-purpose LLMs and led to significant performance gains on extremely difficult real-world SWE-related benchmarks, such as SWE-bench [25]. This raises the question of whether LLM agents could also be useful in automatically setting up and reproducing existing CVEs, allowing us to build reliable, large, and realistic benchmarks for the security community.
In this paper, we present CVE-Genie, a fully-automated, LLM-driven, multi-agent framework for end-to-end CVE reproduction. CVE-Genie instantiates what we define as key properties of an ideal CVE reproduction system, captured by the acronym EAGER. Specifically, an ideal system: Generates working Exploit/proof-of-concept (PoC); Builds Assessors/verifiers for the exploit; Generalizes across CWEs, languages, and project types; Is End-to-end automated; and Rebuilds vulnerable environments. Specifically, this paper makes the following contributions:
@@
    We develop CVE-Genie, an automated end-to-end CVE reproduction pipeline that extracts CVE data (e.g., software vulnerable version source code, advisories, patches, cwes, etc.), rebuilds the vulnerable environment, and generates exploits and verifiers to assess the reproduction.
@@
    We design CVE-Genie to demonstrate compositional intelligence by a novel architecture, systematically selected LLMs and prompts, and reliable EAGER-style CVE reproduction. As a result, CVE-Genie achieves strong performance across diverse CVEs, vulnerability types, programming languages, and projects, even when CVE information is incomplete.
@@
    We systematically evaluate CVE-Genie’s architecture and demonstrate that each component is essential for optimal performance. Notably, even advanced LLMs like o3, when used standalone, are incapable of performing EAGER-style reproduction for even a single CVE.
@@
    We ran CVE-Genie on 841 CVEs (published between June 2024 and May 2025) and successfully reproduced 428 CVEs across 267 projects, 141 CWEs, and 22 programming languages, highlighting our generalizability.
@@
    We make our framework, source code, datasets of reproduced CVEs, and full logs of agent interactions publicly available, providing an ongoing valuable contribution for downstream research tasks (see Appendix C and  F).
@@ 2 Background and Related Work
2.1 Existing Benchmarks and Limitations
To evaluate various vulnerability detection techniques (rule-based methods [44, 58], ML-approaches [36, 30, 31, 29], and LLM-based systems [17, 19, 47, 18]), the community needs standardized benchmarks. Prior work has explored four main approaches for addressing these challenges. (1) Manual efforts involve analyzing [21] or reproducing [66] vulnerabilities in real-world code, or crafting synthetic examples [54], but datasets are small, lack ecological validity, and often suffer labeling errors due to human mistakes (e.g., 50% data in Devign [65] have wrong labels  [50]). (2) Automated labeling using static analysis (SA) tools, such as bandit [6] or infer [23], reduces manual effort but introduces many false positives [48, 64]. (3) Mining real-world data, such as GitHub issues or CVE patch commits, assumes that all removed or modified code is vulnerable; However, this assumption often leads to mislabeled samples, as Risse et al. [50] highlighted for DiverseVul [10] and BigVul [15]. (4) Relying on proven sources, such as OSS-Fuzz, offers verified exploits and sanitizer feedback; However, these are restricted to certain bug types, primarily memory corruption in C/C++, and do not generalize across languages or vulnerability classes [34].
@@ 2.1 Existing Benchmarks and Limitations
2.2 Call for EAGER Style for CVE Reproduction
A CVE ID is a unique identifier for publicly disclosed security vulnerabilities in real-world software. Each CVE ID corresponds to a CVE entry that includes key details for that CVE, such as vulnerability description, source code, affected versions, security advisories, CWE classifications, and patches, enabling security professionals to assess and mitigate risks. Creating a dataset from the standard CVE database is a promising way towards addressing the aforementioned challenges (see Appendix F). Therefore, we introduce EAGER, a set of crucial criteria for ideal CVE reproduction:
@@ 2.2 Call for EAGER Style for CVE Reproduction
However, none of the existing methods satisfy all the criteria (see Table 2). For example, manual effort for producing EAGER datasets does not scale and is not generalizable [38, 66]. ARVO [34] focuses solely on memory-related bugs from OSS-Fuzz C/C++ projects, limiting its scope. In contrast, CVE-Genie is the only framework that satisfies all criteria of EAGER reproduction.
@@
2.3 Large Language Models and Agents
LLMs are transformer-based neural network models with billions of parameters. They have demonstrated remarkable performance across various application domains, including question answering [43], mathematical reasoning [2], and code generation [9, 5]. Recent work also shows their utility in security tasks, e.g., bug detection and repair [46, 57, 24] and writing fuzzer test harnesses [12].
LLM agents are systems that combine LLMs with tools. They can finish complex tasks via a sequence of actions, including planning, generation, and tool executions [53, 22, 45, 27, 26]. An agentic system can have multiple agents, each with a sub-task, such as reasoning, critiquing, generating, perceiving, remembering, or teaching, as well as its own workflow and tool sets. In security, researchers have started to construct LLM agents to find and repair vulnerabilities. There are multi-agent patching systems, where different sub-agents are responsible for fault localization, patch generation, and validation (e.g., PatchPilot [28], and PatchAgent [61]). Through the recent DARPA AIxCC competition, researchers constructed agentic-based systems with end-to-end capabilities for vulnerability detection and patching [18]. These recent successes provide the argument for using LLM agents in vulnerability analysis and software development, as well as for building an EAGER-style automated end-to-end CVE reproduction framework.
@@ 3 CVE-Genie
We design CVE-Genie to exhibit compositional intelligence, i.e., the ability to solve complex tasks by decomposing them into meaningful sub-tasks whose outputs are explicitly composed and verified to form a correct end-to-end solution. To this end, first, we design a novel architecture that decomposes CVE understanding and reproduction into four interdependent modules: (1) the Processor (Section 3.1), which retrieves source code and constructs a structured knowledge base; (2) the Builder (Section 3.2), which reconstructs the vulnerable environment; (3) the Exploiter (Section 3.3), which reproduces the exploit; and (4) the CTF Verifier (Section 3.4), which generates a verifier for the exploit. Second, rather than relying on a single LLM or a fixed prompting strategy across tasks, we systematically select the most suitable LLMs and engineer task-specific prompts for each component of CVE-Genie, resulting in an optimized end-to-end configuration (Section 3.5). Third, we bridge the gap identified in Section 2, namely the lack of any system capable of true EAGER-style CVE reproduction, and ground CVE-Genie’s design in the following guiding principles:
@@
    Modular Task Decomposition: LLMs struggle with long, complex contexts [54] and tasks [25]. Thus, CVE-Genie decomposes reproduction into focused modules and sub-modules handled by specialized agents with systematically engineered prompts, providing the most suitable architecture for end-to-end CVE reproduction (Section 4.2).
@@
    Robustness to Incomplete Data: Since missing CVE fields can lower reproduction success by up to 44% [38], CVE-Genie mitigates this by relying on patch or source-code analysis when advisories/PoCs are unavailable (Section 4.3).
@@
    Reliability through Self-Critique: To counter LLMs’ reasoning limits [54], each module uses paired developer and critic agents in a ReAct-style loop [60], enabling iterative refinement and internal verification (Section 4.2).
Together, these principles elevate CVE-Genie beyond a fixed engineering pipeline and enable CVE-Genie to operationalize compositional intelligence in a fully automated setting, yielding the first practical, generalizable, and end-to-end framework for EAGER-style CVE reproduction (see Appendix E for further details). The following subsections detail the workings of each CVE-Genie component, using CVE-2024-4340 as a running example corresponding to the end-to-end reproduction shown in Figure 2. Moreover, we provide a detailed case study on CVE-2024-5129, a logic vulnerability of missing authorization in lunary-ai, in Appendix H.
Figure 2: CVE-Genie architecture and an end-to-end example of workflow of reproduction for CVE-2024-4340, i.e., Denial of Service due to RecursionError in sqlparse < v0.5.0. See artifact of CVE-2024-4340 complete reproduction run here - https://github.com/BUseclab/cve-genie/tree/main/results/CVE-2024-4340
@@ 3.1 Processor
This module comprises two sub-modules: the Data Processor, which locates the vulnerable project’s source code and gathers relevant raw resources, and the LLM-based Knowledge Builder, which transforms these resources into a structured knowledge base. Below, we provide detailed functionalities of these sub-modules.
@@ 3.1.1 Data Processor 1
For the given CVE ID, this sub-module identifies and extracts the vulnerable version of the software along with all relevant CVE details, as follows:
Source Code Extraction. The Data Processor begins by locating the source code for the vulnerable project linked to the CVE. It searches the GitHub URLs referenced in the “cvelist” 111cvelist” is an automated pilot program for CVE submission through GitHub (https://github.com/CVEProject/cvelist) to identify the project repository. For closed-source software, Data Processor allows users to supply the repository URL. Once the source code is located, it identifies affected software configurations: vulnerable versions, platforms, or settings. For instance, CVE-2024-4340 affects all versions of “sqlparse” prior to v0.5.0. Accordingly, the latest affected version (v0.4.4) is retrieved using the GitHub API, and its source code is downloaded (as shown in Figure 2).
Vulnerability Information Extraction. After the vulnerable version of the project source code is downloaded, Data Processor extracts the following four key pieces of information from “cvelist”, if available: (1) CVE description: A high-level summary of the vulnerability in the target project. (2) CWE data: CWE information, a categorization system for hardware and software weaknesses that associates the CVE with a specific vulnerability type, e.g., CWE-674 (Uncontrolled Recursion). (3) Patch commits: The git-diff of code changes in the target source code that fix the vulnerability. These commits help LLMs localize the issue, understand its root cause [21, 13, 15], and generate effective exploits. (4) Security advisories and PoC: We filter URLs referenced in the CVE entry by keywords (e.g., “security”, “advisory”, “bounty”, etc.), and scrape their contents.
@@ 3.1.2 Knowledge Builder 2
This agent is instructed to analyze, distill, and organize the raw information, extracted by the Data Processor, into a structured knowledge base while retaining essential details required to reproduce the given CVE. This includes extracting PoC or exploit instructions from security advisories, which are invaluable for accurately reproducing the vulnerability. For instance, CVE-2024-4340 has a PoC provided in its security advisory, and the patch commit helps understand the root cause by highlighting where/how the vulnerability was mitigated. As shown in Figure 2, the Knowledge Builder included both of these details in the knowledge base. In case an advisory does not include an exploit, the LLM agent is tasked with generating an overview of what an exploit might entail. Additionally, capturing the patch details helps to localize the root cause of the vulnerability [21, 13, 7] and aids in crafting an effective exploit. This knowledge base acts as the long-term memory for agents, helping them to effectively reproduce CVEs, by including only essential information in their context to avoid overloading the LLM’s context window.
Output. A structured CVE knowledge base providing details of CVE, including its associated CWEs, affected software configurations, root causes (if provided in the security advisory), and essential information from security advisories and patch commits.